A body falling from a vertical height of 10m pierces through a distance of 1m in sand. It faces an average
retardation in sand equal to?

[Answer is in terms of g i.e acceleration due to gravity ]

Grade:

2 Answers

shubham sharma
10 years ago

Ans is 10g.

Just find vel. With which it strikes down and then for rest motion apply equations of motions.

Nikita Gidwani
6 years ago
u=0g=10Velocity just before piercing in sand v^2-0=2*10*9v^2=180m/sRetardation experienced in sand 0-180=2*(-a)*1a=90=9*10=9gHence retardation=9g